myininaya (myininaya):

integrate e^(sqrt(3x+9))

myininaya (myininaya):

\[\int\limits_{}^{}e^{\sqrt{3x+9}}dx\]

OpenStudy (anonymous):

cant

OpenStudy (anonymous):

huh , wait, wolframa says you can lols

myininaya (myininaya):

try let u=sqrt(3x+9) play with that

OpenStudy (anonymous):

u= 3x+9 and use int by parts hmm havent seen that before

myininaya (myininaya):

u^2=3x+9 2du=3dx and right elecegineer!

myininaya (myininaya):

oops 2udu=3dx

OpenStudy (anonymous):

e^sqrt3x+9.3/2(sqrt3x+9)

OpenStudy (anonymous):